The sum of the ages of Erik and Lauren is 111 years. 8 years ago, Erik's age was 4 times Lauren's age. How old is Erik now?
Answer by ikleyn(52794)

.

    8 years ago, Eric's age was 4 times Lauren's age, AND

AND 8 years ago, the sum of their ages was  111 - 8 - 8 = 95.



If x is the Lauren's age 8 years ago, the the Eric's age was  4x,

and you have this equation


    x + 4x = 95

    5x     = 95

     x     = 95/5 = 19.


So, 8 years ago, Lauren was 19 years old;  Eric was  19*4 = 76 years old.


Hence, now their age is  19+8 = 27  and  76+8 = 84 years, respectively.    ANSWER

Solved.